OpenStudy (anonymous):

5y^8-125 Factor completely I need help doing this !

OpenStudy (id21):

\[5y^8-125=5y^8-5^3=5(y^8-25)\]

OpenStudy (anonymous):

\[=5(y^8-25)\] \[=5(y^4-5)(y^4+5)\]